OpenStudy (anonymous):

find a so that the point (-1, 2) is on the graph of f(x)=ax^2+5

hartnn (hartnn):

since (-1,2) is on the graph f(x)=ax^2+5 you can put f(x)=2, and x=-1 what u get ?

OpenStudy (anonymous):

2=a(-1)+5

hartnn (hartnn):

2=a(-1)^2+5 find a from here

hartnn (hartnn):

don't miss ^2

OpenStudy (anonymous):

a = ..

hartnn (hartnn):

... ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

?

hartnn (hartnn):

very close,but not correct 2=a+5 a=?

OpenStudy (anonymous):

-3

hartnn (hartnn):

that is correct value of a

OpenStudy (anonymous):

thank you again i also have a question just like the one we did a question ago. do you have time?

hartnn (hartnn):

few minutes only...

OpenStudy (anonymous):

4x^2 + 25y^2 = 100

hartnn (hartnn):

intercepts?

OpenStudy (anonymous):

yes there will be four i know.

OpenStudy (anonymous):

25,0 -25,0 for x?

hartnn (hartnn):

nopes

OpenStudy (anonymous):

y 0 -4 0 4

hartnn (hartnn):

nopes

OpenStudy (anonymous):

darn.. can u help?

hartnn (hartnn):

u got x^2 =25 right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

those will be x ints? (-5,0) (5,0)

hartnn (hartnn):

\(\checkmark\)

OpenStudy (anonymous):

0,-2 0,2

OpenStudy (anonymous):

y

hartnn (hartnn):

that is also \(\checkmark\)

OpenStudy (anonymous):

thank u so much. when r u on next? u should be my reg helper.

hartnn (hartnn):

idk, i come online when i m free...

OpenStudy (anonymous):

oh ok.